A motorboat can maintain a constant speed of 16 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 20 minutes; the return trip takes 15 minutes. What is the speed of the current?
Answer by mananth(16949)

speed of current be x mph
let the distance traveled be y
upstream speed = 16-x
downstream speed = 16+x mph
distance / speed = time
y/(16-x)= 1/3
3y=(16-x)
3y=16-x
3y+x=16-------------------(1)
y/(16+x)= 1/4
4y=(16+x)
4y=16+x
4y-x=16...................(2)
x+3y=16 .............1
-x+4y= 16 .............2
Eliminate
multiply (1)by -4
Multiply (2) by 3
-4x-12y =-64
-3x+12y =48
Add the two equations
-7x=-16
/ -7
x=2.29
plug value of x in (1)
x+3y=16
2.29+3y =16
3y = 16-2.29
3y = 13.71
y = 4.57 miles
speed of current = 2.29 mph